OpenStudy (anonymous):

If two fair dice are rolled, what is the probability that a total of four shows?

OpenStudy (anonymous):

there are 3 ways by which we will get 4 as the sum. So, 3/36 = 1/12 is the answer.

OpenStudy (kropot72):

The sample space has 36 possible outcomes. E = {(3, 1), (2, 2), (1, 3)}. These are the only combinations of numbers on the two dice which give a sum of four. P(E) = 3/36 = 1/12